OpenStudy (anonymous):

Find the parabola of the form y=ax^2+b which best fits the points (1,0), (5,5), (7,10) by minimizing the sum of squares, S, given by S=(a+b)^2+(25a+b−5)^2+(49a+b−10)^2

OpenStudy (goformit100):



OpenStudy (anonymous):

thank you both

OpenStudy (anonymous):

solve:\[0=a+b\\ 5=25a+b 24a=5\implies a=5/24 b=-5/24 \]

OpenStudy (anonymous):

Thank you